OpenStudy (mendicant_bias):

I'm trying to use s-axis translation properties to evaluate an inverse laplace transform, and I got a solution, but I'm missing part of the book's soln and can't figure out why; could somebody take a look at this?

OpenStudy (mendicant_bias):

\[L^{-1} \left\{\vphantom{} \frac {s}{s^2+4s+5} \right\}\]

OpenStudy (mendicant_bias):

\[L^{-1} \left\{\vphantom{} \frac{s}{s^2+4s+5} \right\} = L^{-1} \left\{\vphantom{} \frac {s}{(s+2)^2+1} \right\}\]

OpenStudy (mendicant_bias):

Fitting the general form of an inverse laplace transform of cosine(kt), \[L^{-1} \left\{\vphantom{} \frac {s}{s^2 + k^2} \right\}\]...whoop, wait, it's right there, I see it, nevermind. The numerator has to be subtracted something to fit that same form of \[s \pm c\] in order to be dealt with in that way? Wait, maybe?

OpenStudy (mendicant_bias):

\[L^{-1} \ Y(s) = L^{-1} \left\{\vphantom{} \frac{s}{(s+2)^2+1} \right\};\] \[L^{-1} Y(s) +L^{-1} \left\{\vphantom{} \frac{2}{s^2+4s+5} \right\} = L^{-1} \left\{\vphantom{} \frac{s}{s^2+4s+5} \right\} +L^{-1} \left\{\vphantom{} \frac{2}{s^2+4s+5} \right\}\] \[L^{-1} Y(s) +L^{-1} \left\{\vphantom{} \frac{2}{s^2+4s+5} \right\} =L^{-1} \left\{\vphantom{} \frac{s+2}{s^2+4s+5} \right\} \] \[L^{-1} Y(s)=L^{-1} \left\{\vphantom{} \frac{s+2}{s^2+4s+5} \right\} -L^{-1}\left\{\vphantom{} \frac{2}{s^2+4s+5} \right\} \]

OpenStudy (anonymous):

Okay so far. Next step would be to use the completed-square form of each denominator.

OpenStudy (anonymous):

\[F(s+2)=\frac{s+2}{(s+2)^2+1}~~\implies~~F(s)=\frac{s}{s^2+1}\] Use the fact that \[\mathcal{L}^{-1}\{F(s-c)\}=e^{-cs}f(t)\]

OpenStudy (mendicant_bias):

Yep, taking a shot now:

OpenStudy (mendicant_bias):

http://i.imgur.com/XmHIdo1.png This look right?

OpenStudy (anonymous):

Yes that's correct.